Sharon and Rae are riding bicycles on the same trail in the same direction, shareon crosses a bridge at 12:18 pm, Rae crosses the same bridge at 12:33pm, If sharon is going 8mph and Rae is going 10mph, when will Rae catch up with sharon?
Found 2 solutions by ankor@dixie-net.com, Maths68:
Answer by ankor@dixie-net.com(22740)

Sharon and Rae are riding bicycles on the same trail in the same direction, shareon crosses a bridge at 12:18 pm, Rae crosses the same bridge at 12:33pm,
If sharon is going 8mph and Rae is going 10mph, when will Rae catch up with sharon?
:
From the information given, we know that R is 15 min behind S, at 12:18 PM
convert 15 min to .25 hrs
Find the distance R is behind S at this time: .25(10) = 2.5 mi
:
let t = time in hrs, that R catches s
:
Write a distance equation; dist = speed * time
:
10t = 8t + 2.5
10t - 8t = 2.5
2t = 2.5
t = 2.5/2
t = 1.25 hrs, which is 1 hr 15 min
:
Find the time from 12:18 PM: 12:18 + 1:15 = 13:33, or 1:33 PM, R catches S
:
:
Check this by finding the distance each rode
10*1.25 = 12.5 mi distance traveled
8*1.25 = 10 mi + 2.5 mi head start: = 12.5

Answer by Maths68(1474)

Sharon Data
===========
Distance traveled = x
Speed = 8mph
Time taken to cover distance x = t
Distance = Speed * time
x=8*t
Rae Data
========
Distance traveled = x
Speed = 10mph
Time taken to cover distance x = t-15 (Rae started 15 minutes later)
Distance = Speed * time
x=10*(t-15)
Since both will cover the same distance therefore
10(t-15)=8t
10t-150=8t
10t-8t=150
2t=150
t=150/2
t=75 minutes or 1.25 hours
After (75-15) 60 minutes of crossing the bridge Rea will catch up with Sharon at 1:33pm
Check
=======
Sharon Data
Distance traveled = x
Speed = 8mph
Time taken to cover distance x = t = 1.25 hours
Distance = Speed * time
x=8*t= 8*1.25 = 10 miles
Rae Data
Distance traveled = x
Speed = 10mph
Time taken to cover distance x = t-15 = 75-15=60 minutes = 1 hour
Distance = Speed * time
x=10*1= 10 miles